OpenStudy (anonymous):

A horse canters away from its trainer in a straight line, moving 48m away in 9.1s . It then turns abruptly and gallops halfway back in 2.0s .Calculate its average speed for the entire trip

OpenStudy (anonymous):

total distance covered = 48+48 = 96 total time taken = 9.1+2=11.2 therefore average speed = 96/11.2=8.572